### v2

# Elastic-plastic cell-based smoothed finite element method solving geotechnical problems

Abstract
An elastic–plastic cell-based smoothed finite element method (CSFEM) is proposed for geotechnical analysis of soils and rocks exhibiting nonlinear and path-dependent behaviors. By introducing strain smoothing over subcell domains and employing a consistent stress return-mapping algorithm, the method enhances stress accuracy, alleviates volumetric locking, and reduces sensitivity to mesh distortion while retaining the flexibility of polygonal elements. The formulation is implemented in ABAQUS via a user-defined element and validated through benchmark and practical problems, including a pressurized thick cylinder, biaxial soil test, strip footing bearing capacity, tunnel excavation, and slope stability. Numerical results show excellent agreement with analytical solutions and conventional FEM, with smoother stress fields, improved convergence, and higher accuracy in ultimate load prediction. These findings demonstrate that CSFEM provides a stable and efficient framework for elastic–plastic analysis of complex geotechnical problems.
keywords:
cell-based smoothed finite element method (CSFEM), elastic–plastic analysis, geotechnical engineering, ABAQUS UEL implementation
[…]
[inst1]organization=Department One,addressline=Address One, city=City One, postcode=00000, state=State One, country=Country One
[…]
[inst2]organization=Department Two,addressline=Address Two, city=City Two, postcode=22222, state=State Two, country=Country Two
1 Introduction
The reliable prediction of the mechanical response of geomaterials under complex loading conditions remains a fundamental challenge in geotechnical engineering. Geomaterials, such as soil and rock, often exhibit nonlinear behaviors including plasticity, strain-softening, and path-dependent responses (Chu et al., 1992; Kok et al., 2009). These features require robust numerical methods to capture the intricate interplay between elasticity and plasticity in large-scale simulations of foundations (Ray et al., 2021), slopes (Kardani et al., 2021), tunnels (Jin et al., 2022), embankments(Xia et al., 2024), and underground structures(Hemeda, 2022). Among various computational approaches, the finite element method (FEM) has become the most widely adopted due to its flexibility in handling irregular geometries and boundary conditions (Bottero et al., 1980; Liu et al., 2015; Wei et al., 2025).
However, FEM formulations present several drawbacks in the elastic–plastic analysis of geotechnical problems. FEM is known to overestimate the stiffness matrix, leading to an overall response that is too stiff and an underestimation of the internal strain energy (Fraeijs de Veubeke, 1965). As a result, the computed displacements are generally smaller than the actual values. In addition, the required mapping and coordinate transformation restrict element shapes: for a four-node isoparametric element, all interior angles must remain below 180 and the Jacobian determinant must remain positive (Liu and Quek, ; Liu et al., ; Liu et al., 2007). These constraints increase computational cost and significantly limit FEM performance in geotechnical problems involving large deformation or severe mesh distortion.
By introducing the strain-smoothing technique first proposed by Liu et al. (Liu et al., 2007), the smoothed finite element method (SFEM) (Liu and Trung, 2016) has developed into a promising framework for overcoming the intrinsic limitations of classical FEM, thereby drawing substantial attention in recent years. The SFEM delivers higher accuracy and faster energy convergence than the conventional FEM (Dai et al., 2007; Yan et al., 2025). Specifically, its freedom from mapping or coordinate transformation allows elements of arbitrary shapes, making it particularly beneficial for geotechnical analyses where large deformations can cause severe mesh distortion.
Several SFEM variants have been proposed, including face-based (Huang et al., 2023; Jiang et al., 2025), node-based (Sun et al., 2022; Lyu et al., 2024), edge-based (He, 2022; He et al., 2022), and cell-based formulations (Surendran et al., 2021; Liu et al., 2022; Sun et al., 2025).Among these, the cell-based SFEM (CSFEM) provides a particularly straightforward framework that remains close to the standard FEM procedure. In the CSFEM, each element is subdivided into smoothing sub-cells where the strain field is averaged, and the stiffness contribution is evaluated through surface integration based on the divergence theorem (Lee et al., 2015).
At present, the CSFEM has been successfully applied to several physical field problems: Surendran et al.(Surendran et al., 2021) presents a CSFEM using polygonal elements to model interfacial cracks with non-matching grids, enabling accurate fracture analysis while reducing degrees of freedom and relaxing meshing constraints. Yan et al. (Yan et al., 2025) present a fast CSFEM to solve static and dynamic problems. Lee et al. (Lee et al., 2015) a three-dimensional CSFEM for application to elastic-plastic analysis for the solid mechanics. Peng et al. (Peng et al., 2021) developed a novel phase-field model for brittle fracture using CSFEM with spectral decomposition to capture nonlinear stress and elastic responses. These studies demonstrate the effectiveness and versatility of CSFEM, yet its application in geotechnical engineering simulations remains limited.
In this work, we propose an elastic–plastic CSFEM tailored for geotechnical applications. The method incorporates smoothing operations into standard FEM formulations. The proposed framework enhances numerical stability, reduces sensitivity to mesh distortion, and improves the accuracy of stress prediction in nonlinear problems. The remainder of this paper is organized as follows. Section 2 outlines the theoretical framework of the proposed elastic-plastic CSFEM, including the smoothing formulation and elastic–plastic integration scheme. Section 5 details the numerical implementation and integration of elastic-plastic CSFEM into ABAQUS via a user-defined element (UEL). Section 6 presents benchmark numerical studies to verify accuracy and convergence. Finally, Section 7 concludes with key findings and future research directions.
2 Cell-based SFEM formulation for elasto-plasticity
In this study, we examine a deformable body that occupies the domain $\Omega$. This body responds to internal body forces b, externally applied tractions $\overline{\textbf{t}}$ on the boundary $\Gamma_{t}$, and displacement boundary conditions $\mathbf{u}=\overline{\mathbf{u}}$ on $\Gamma_{u}$, as depicted in Figure 1. The governing equation and boundary conditions are expressed as follows:
$\begin{array}[]{rlc}\nabla\boldsymbol{\sigma}+\mathbf{b}=0&\text{ in }&\Omega\\ \mathbf{u}=\overline{\mathbf{u}}&\text{ in }&\Gamma_{u},\\ \boldsymbol{\sigma}\mathbf{n}=\overline{\mathbf{t}}&\text{ in }&\Gamma_{t}\end{array}$ (1)
where $\nabla$ denotes the differential operator, $\boldsymbol{\sigma}$ represents the Cauchy stress tensor.
Figure 1: A two-dimensional deformable body.
The relationship of stress $\boldsymbol{\sigma}$ and strain $\boldsymbol{\varepsilon}$ can be expressed by
$\boldsymbol{\sigma}=\boldsymbol{D}\boldsymbol{\varepsilon},$ (2)
where $\boldsymbol{D}$ is the constitutive matrix.
Mathematically expressing the virtual displacement and compatible strain within each element of the spatial discretization, we obtain the subsequent equations:
$\delta\mathbf{u}^{h}=\sum_{i}\mathbf{N}_{i}\delta\mathbf{d}_{i},\quad\mathbf{u}^{h}=\sum_{i}\mathbf{N}_{i}\mathbf{d}_{i},$ (3)
$\delta\boldsymbol{\varepsilon}^{h}=\sum_{i}\mathbf{B}_{i}\delta\mathbf{d}_{i},\quad\boldsymbol{\varepsilon}^{h}=\sum_{i}\mathbf{B}_{i}\mathbf{d}_{i},$ (4)
where $\mathbf{d}_{i}=\left[\begin{array}[]{ll}u_{i}&v_{i}\end{array}\right]^{\mathrm{T}}$ is the nodal displacement vector, and $\boldsymbol{B}_{i}$ is the strain-displacement matrix as follows:
$\mathbf{B}_{i}=\left[\begin{array}[]{cc}N_{ix}&0\\ 0&N_{iy}\\ N_{iy}&N_{ix}\end{array}\right],$ (5)
where $N_{ix}$ and $N_{iy}$ are outer normal derivatives in relation to $x$ and $y$. The process of energy assembly yields the following results:
$\int_{\Omega}\delta\mathbf{d}^{\mathrm{T}}\mathbf{B}^{\mathrm{T}}\boldsymbol{\sigma}\mathrm{d}\Omega-\int_{\Omega}\delta\mathbf{d}^{\mathrm{T}}\mathbf{N}^{\mathrm{T}}[\mathbf{b}-\rho\ddot{\mathbf{u}}-c\mathbf{u}]\mathrm{d}\Omega-\int_{\Gamma_{t}}\delta\mathbf{d}^{\mathrm{T}}\mathbf{N}^{\mathrm{T}}\mathbf{t}\mathrm{d}\Gamma=0,$ (6)
As the provided expression remains valid for any arbitrary virtual displacement , we can conclude that:
$\int_{\Omega}\mathbf{B}^{\mathrm{T}}\boldsymbol{\sigma}\mathrm{d}\Omega-\int_{\Omega}\mathbf{N}^{\mathrm{T}}\mathbf{b}\mathrm{d}\Omega-\int_{\Gamma_{t}}\mathbf{N}^{\mathrm{T}}\mathbf{t}\mathrm{d}\Gamma=0.$ (7)
Following this, we can express the discrete governing equation as following:
$\mathbf{Kd}=\mathbf{P}.$ (8)
$\mathbf{K}=\int_{\Omega}\mathbf{B}^{T}\mathbf{DB}\mathrm{d}\Omega,$ (9)
$\mathbf{P}=\int_{\Omega}\mathbf{N}^{T}\mathbf{b}\mathrm{d}\Omega+\int_{\Gamma_{t}}\mathbf{N}^{T}\mathbf{t}\mathrm{d}\Gamma,$ (10)
where $\mathbf{K}$ is the stiffness, mass, and damping matrices respectively; $\mathbf{P}$ is the external force vector.
The modified strain field $\tilde{\boldsymbol{\varepsilon}}^{h}$ is determined through a weighted-average formulation of the conventional strain field
$\tilde{\boldsymbol{\varepsilon}}^{h}=\int_{\Omega_{C}}\boldsymbol{\varepsilon}^{h}(\mathbf{x})\Phi\left(\mathbf{x}-\mathbf{x}_{C}\right)\mathrm{d}\Omega,$ (11)
where $\Omega_{C}$ denotes the sub-cell, and $\Phi\left(\mathbf{x}-\mathbf{x}_{C}\right)$ represents the smoothing function, which is defined as:
$\Phi\left(\mathbf{x}-\mathbf{x}_{C}\right)=\begin{cases}1/A_{C}&\mathbf{x}\in\Omega_{C}\\ 0&\mathbf{x}\notin\Omega_{C}\end{cases},$ (12)
where $A_{C}$ is the area of the sub-cell. The modified strain field $\tilde{\boldsymbol{\varepsilon}}^{h}$ is computed from the smoothed discretized gradient operator $\widetilde{\mathbf{B}}$ and the nodal displacement $\mathbf{b}$. By applying the technique of integration by parts to Eq.(11), we obtain:
$\tilde{\varepsilon}^{h}\left(\mathbf{x}_{C}\right)=\int_{\Gamma}\mathbf{u}^{h}(\mathbf{x})\mathrm{n}(\mathbf{x})\Phi\left(\mathbf{x}-\mathbf{x}_{C}\right)d\Gamma-\int_{\Omega}\mathbf{u}^{h}(\mathbf{x})\nabla\Phi\left(\mathbf{x}-\mathbf{x}_{C}\right)d\Omega.$ (13)
Upon substituting Eq.(12) into Eq.(13), the smoothed gradient or strain field can be derived
$\tilde{\boldsymbol{\varepsilon}}^{h}\left(\mathbf{x}_{C}\right)=\int_{\Gamma_{C}}\mathbf{u}^{h}(x)\mathbf{n}(x)\Phi\left(\mathbf{x}-\mathbf{x}_{C}\right)d\Gamma=\frac{1}{A_{c}}\int_{\Gamma_{c}}\mathbf{u}^{h}(\mathbf{x})\mathbf{n}(\mathbf{x})d\Gamma,$ (14)
where $\Gamma_{C}$ represents the boundary of the sub-cell. The smoothed strain can be expressed as:
$\tilde{\boldsymbol{\varepsilon}}\left(\mathbf{x}_{C}\right)=\sum_{I}^{n}\widetilde{\mathbf{B}}_{I}\left(\mathbf{x}_{C}\right)\mathbf{d}_{I},$ (15)
where $n$ denotes the number of element nodes. The stiffness matrix computed for the CSFEM is presented as follows:
$\widetilde{\mathbf{K}}=\sum_{I=1}^{nel}\widetilde{\mathbf{K}}^{h}=\sum_{I=1}^{nel}\int_{\Omega^{e}}\widetilde{\mathbf{B}}^{\mathrm{T}}\mathbf{D}\widetilde{\mathbf{B}}\mathrm{d}\Omega.$ (16)
The derivation process also reveals that the strain-displacement matrix
$\overset{\sim}{\mathbf{B}}=\begin{bmatrix}\overset{\sim}{\mathbf{B}}_{1}&\overset{\sim}{\mathbf{B}}_{2}&\cdots&\overset{\sim}{\mathbf{B}}_{n}\end{bmatrix}$, (17)
where the strain-displacement matrix can be written as
$\overset{\sim}{\mathbf{B}}_{k}=\frac{1}{A_{c}}\int_{\Gamma_{c}^{h}}\begin{bmatrix}n_{x}(\mathbf{x})&0\\ 0&n_{y}(\mathbf{x})\\ n_{y}(\mathbf{x})&n_{x}(\mathbf{x})\end{bmatrix}N_{k}(\mathbf{x})\mathrm{d}\Gamma.$ (18)
3 Elasto-plastic constitutive formulation
In elasto-plastic analysis, the total incremental strain is conventionally decomposed into its elastic and plastic components as
$\boldsymbol{\Delta\varepsilon}=\boldsymbol{\Delta\varepsilon}^{\mathrm{e}}+\boldsymbol{\Delta\varepsilon}^{\mathrm{p}},$ (19)
where $\boldsymbol{\Delta\varepsilon}^{\mathrm{e}}$ and $\boldsymbol{\Delta\varepsilon}^{\mathrm{p}}$ denote the elastic and plastic strain increments, respectively. The incremental stress–strain relationship for the elastic response follows Hooke’s law, given by
$\boldsymbol{\Delta\sigma}=\mathbf{D}:\left(\boldsymbol{\Delta\varepsilon}-\boldsymbol{\Delta\varepsilon}^{\mathrm{p}}\right)=\mathbf{D}:\boldsymbol{\Delta\varepsilon}^{\mathrm{e}},$ (20)
where $\mathbf{D}$ represents the elastic constitutive matrix.
The evolution direction of plastic strain is governed by the plastic flow rule. For associative plasticity, the plastic strain increment is determined from the yield function $F$ and the plastic multiplier $\Delta\lambda$ as
$\boldsymbol{\Delta\varepsilon}^{\mathrm{p}}=\frac{\partial g}{\partial\boldsymbol{\sigma}}\,\Delta\lambda=\frac{\partial F}{\partial\boldsymbol{\sigma}}\,\Delta\lambda,$ (21)
where $g$ is the plastic potential function, which equals the yield function $F$ in the associative case. When the stress state reaches the yield surface, the consistency condition must be satisfied to maintain stress continuity, expressed as
$\Delta F=\frac{\partial F}{\partial\boldsymbol{\sigma}}:\boldsymbol{\Delta\sigma}+\frac{\partial F}{\partial\kappa}\,\Delta\kappa=0,$ (22)
where $\kappa$ denotes the hardening parameter. The evolution of $\kappa$ is described by the hardening law
$\Delta\kappa=H\,\Delta\lambda,$ (23)
with $H$ being the plastic hardening modulus.
In this study, the Mohr–Coulomb failure criterion is adopted to assess the stability of vegetated slopes. The yield function is defined as
$F=\left(\sqrt{3}\cos\theta_{0}-\sin\theta_{0}\sin\phi\right)q-3p\sin\phi-3(c_{r}+c_{s})\cos\phi=0,$ (24)
where $p=\tfrac{1}{3}(\sigma_{1}+\sigma_{2}+\sigma_{3})$ is the mean stress, $q=\sqrt{\tfrac{3}{2}\boldsymbol{s}:\boldsymbol{s}}$ is the equivalent deviatoric stress, $\theta_{0}$ is the Lode angle, $\phi$ is the internal friction angle, and $c_{r}$ and $c_{s}$ denote the root and soil cohesion, respectively.
By combining the flow rule, the consistency condition, and the hardening law, the elasto-plastic constitutive matrix can be derived as
$\mathbf{D}_{ep}=\mathbf{D}-\mathbf{D}\left(\frac{\partial F}{\partial\boldsymbol{\sigma}}\right)\left(\frac{\partial F}{\partial\boldsymbol{\sigma}}\right)^{T}\mathbf{D}\left(A+\left(\frac{\partial F}{\partial\boldsymbol{\sigma}}\right)^{T}\mathbf{D}\left(\frac{\partial F}{\partial\boldsymbol{\sigma}}\right)\right)^{-1},$ (25)
where $A=-\dfrac{1}{\Delta\lambda}\dfrac{\partial F}{\partial\kappa}\Delta\kappa$ and $\Delta\lambda$ is the plastic multiplier. The elasto-plastic constitutive matrix $\mathbf{D}_{ep}$ provides the incremental stress–strain relationship during yielding and varies spatially within the domain, reflecting the nonlinear mechanical response of the material.
4 Construction of CSFEM elements
In the CSFEM, each element is subdivided into several smoothing domains to construct the stiffness matrix. As shown in Fig. 2, a quadrilateral element is divided into four subcells, within which the strain field is smoothed and the domain integration is transformed into boundary integrals. Since only the shape function values along the boundaries are required, simple interpolation schemes are used at the Gauss points located on the element edges. Representative shape function values at typical points within the element are listed in Tab. 1, which are used to evaluate the smoothed strain field and the corresponding stiffness matrix.
Figure 2: Construction of a cell-based smoothing domain.
Table 1: Shape function value at different sites within an element (Fig. 2)
Site Node 1 Node 2 Node 3 Node 4 Description
1 1.0 0 0 0 Field node
2 0 1.0 0 0 Field node
3 0 0 1.0 0 Field node
4 0 0 0 1.0 Field node
5 0.5 0.5 0 0 Side midpoint
6 0 0.5 0.5 0 Side midpoint
7 0 0 0.5 0.5 Side midpoint
8 0.5 0 0 0.5 Side midpoint
9 0.25 0.25 0.25 0.25 Centroid
The use of smoothed strain fields in CSFEM leads to a softer stiffness matrix compared with the conventional FEM, which helps alleviate numerical issues such as mesh distortion sensitivity and volumetric locking. Furthermore, the method retains simplicity in implementation since only the original finite element shape functions are required at the boundary integration points, without computing their spatial derivatives.
5 Implementation
5.1 Implementation of the elastic-plastic CSFEM
Fig. 3 and Algorithm 1 illustrate the computational workflow of the elastic–plastic analysis implemented in ABAQUS via the UEL interface. The main responsibility of the UEL subroutine is to evaluate and update the element contributions to the internal force vector (RHS) and stiffness matrix (AMATRX), based on the information supplied by the ABAQUS/Standard solver (Dassault Systèmes Simulia Corp., 2010). Within this framework, the stiffness matrix and the residual force vector are formulated as
$\mathrm{AMATRX}=\tilde{\mathrm{K}}$, (26)
$\mathrm{RHS}=-\tilde{\mathrm{KU}}_{t+\Delta t}.$ (27)
where $\tilde{\mathrm{K}}$ denotes the element stiffness matrix and $\mathrm{U}_{t+\Delta t}$ represents the displacement vector at the end of the current increment. The UEL routine iteratively updates both AMATRX and RHS according to Eqs. (26) and (27) throughout the solution process.
Figure 3: Framework sketch of implementing the elastic-plastic CSFEM within ABAQUS
Algorithm 1 Solving the CSFEM UEL
1:Node and element information, material properties, nodal displacement $u_{t}$
2:Nodal displacement $u_{t+1}$
3:Iteration: increment steps $k=1$
4:while ABAQUS not converged do
5:  solve nodal displacement $u^{k}_{t+1}$
6:  for $i=1$ to $n_{c}$ do
7:   solve shape functions $N_{k}(x)$ in Table 1
8:   solve strain–displacement matrix $\tilde{\mathrm{B}}_{c}$ in Eq.(18)
9:   solve elasto-plastic constitutive matrix $\mathrm{D}_{ep}$ in Eq.(25)
10:   solve stiffness matrix of subcells $\tilde{\mathrm{K}}_{c}$ in Eq.(16)
11:  end for
12:  obtain element matrices $\tilde{\mathrm{K}}$
13:  update AMATRX and RHS according to Eqs. (26) and (27)
14:  update $k=k+1$
15:end while
16:solve nodal field variables $u_{t+1}=u^{k}_{t+1}$
5.2 Visualization of results
As shown in Fig. 4(a), the UEL in ABAQUS is represented by an “×” symbol, and its stress–strain results cannot be visualized directly. To address this limitation, a data-linking strategy was developed between the UEL and the user material subroutine (UMAT), as schematically illustrated in Fig. 3. Through this interface, the stress, strain, and other state variables (STATEV) computed within the UEL were transferred to the UMAT framework, allowing the use of standard ABAQUS elements for post-processing and visualization.
For visualization in ABAQUS/CAE, these quantities were exported via state variables declared in the input file using the *DEPVAR option and were stored at the integration points during the analysis. Consequently, as shown in Figs. 4(b) and (d), the numerical results obtained from the UEL could be accessed and visualized within the standard post-processing environment, providing clear field distributions comparable to those produced by conventional built-in elements.
Figure 4: Visualization of results
6 Numerical examples
6.1 Pressurized thick-cylinder modeled using a quarter-annulus model
In this example, we considered a quarter-thickness cylindrical model subjected to internal pressure to verify the proposed method. The displacements in the polar coordinate system can be derived analytically (Timoshenko and Goodier, 1970).
${U}_{\mathrm{rad}}=\frac{R_{a}^{2}Pr}{E\left(R_{b}^{2}-R_{a}^{2}\right)}\left(1-\nu+\left(\frac{R_{b}}{r}\right)^{2}(1+\nu)\right),$ (28)
$u_{x}=u_{\mathrm{rad}}\cos\theta,$ (29)
$u_{y}=u_{\mathrm{rad}}\sin\theta.$ (30)
where $R_{a}$ and $R_{b}$ denote the inner and outer radii of the cylinder, respectively, with $R_{a}=1\,\text{m}$ and $R_{b}=2\,\text{m}$. A uniform pressure of $1000\,\text{Pa}$ was applied along the inner circular boundary, while the material parameters were defined by Young’s modulus $E=10000$ Pa and Poisson’s ratio $\nu=0.25$. The geometry of the cylinder and the associated boundary conditions are illustrated in Fig. 5. The meshes are successively refined to element sizes of 0.125 m, 0.0625 m, 0.03 m, and 0.015 m, as shown in Fig. 6.
Figure 5: Fixed constraint beam mesh model at both ends
Figure 6: Four levels of mesh refinement; (a) 0.125 m mesh size; (b) 0.0625 m mesh size; (c) 0.03 m mesh size; (d) 0.015 m mesh size.
Figure 7: Convergence of the relative error in the displacement.
Figure 8: Contour plots of results in thick-cylinder; (a) the FEM; (b) the CSFEM.
6.2 Biaxial test
In this example, we consider a typical biaxial test setup used to investigate the mechanical behavior of soils, as shown in Fig. 9 (a). The soil specimen is initially subjected to prescribed confining stresses of 100 kPa in both the vertical and horizontal directions. After the initial stress state is established, a vertical displacement is incrementally imposed while the lateral stress is kept constant until failure occurs. The soil is modeled with an elastic modulus $E=10$ MPa, Poisson’s ratio $\nu=0.3$, cohesion $c=10$ kPa, and an internal friction angle of $30^{\circ}$.
For the present biaxial test, adopting the Mohr-Coulomb failure criterion with cohesion $c$ and friction angle $\varphi$, and taking the major and minor principal stresses as $\sigma_{1}$ and $\sigma_{3}$, respectively, the ultimate major principal stress is given by
$\sigma_{1,\mathrm{ult}}\;=\;\sigma_{3}\;\frac{1+\sin\varphi}{1-\sin\varphi}\;+\;\frac{2c\cos\varphi}{1-\sin\varphi}.$ (31)
According to Eq. (31), the ultimate major principal stress in this example is 334.64 kPa. Fig. 10 shows the relationship between vertical stress and vertical strain. As illustrated in Fig. 10, during the initial loading stage, the material remains in the elastic region, and the vertical stress increases with the increase in vertical strain until yielding occurs. Fig. 11 illustrates the vertical stress and displacement contours at the material yielding. It can be observed from Fig. 11 that the final vertical stress obtained by CSFEM and FEM is 334.6 kPa, which agrees well with the result of Eq. (31). Moreover, the displacement distributions of CSFEM and FEM are also essentially consistent.
Figure 9: Schematic diagram for the bearing capacity problem; (a) geometry and boundary conditions; (b) mesh model.
Figure 10: The relationship between vertical stress and vertical strain.
Figure 11: The vertical stress and displacement contours at the material yielding.
6.3 Bearing capacity problem
In this example, we consider a flexible strip footing resting on the surface of a half space subjected to a uniform vertical load, as shown in Fig. 8(a). Due to symmetry, only half of the footing is modeled. Roller boundary conditions are imposed on the side edges, while the bottom boundary is fully fixed. The analysis is carried out under plane strain conditions using the mesh shown in Fig. 8(b), with 80 load increments of 8000 Pa. The corresponding soil properties are given in Table 2. For a soil with zero unit weight ($\gamma=0$), the ultimate bearing capacity of strip foundations can be expressed as:
$q_{u}=\begin{cases}c\cdot N_{c}&\phi\neq 0\\ (\pi+2)\cdot c&\phi=0\end{cases}\quad,$ (32)
where $N_{c}$ is a bearing capacity factor can be given by,
$N_{c}=(N_{q}-1)\cot\phi,$ (33)
$N_{q}=e^{\pi\tan\phi}\left(\frac{1+\sin\phi}{1-\sin\phi}\right).$ (34)
Table 2: Soil parameters of a flexible strip footing
$E$ (kPa) $\nu$ $\phi$ $c$ (kPa) $\psi$
10000 0.3 $5^{\circ}$ 1 $5^{\circ}$
[…]
Figure 12: Schematic diagram for the bearing capacity problem; (a) geometry and boundary conditions; (b) mesh model.
Figure 13: Bearing capacity–displacement diagram of a flexible footing using CSFEM and FEM.
Table 3: Relative errors in bearing capacity between CSFEM and FEM
Analytical solution (Pa) CSFEM (Pa) FEM (Pa) Relative error of CSFEM Relative error of FEM
[…]
Figure 14: Comparison of vertical displacement and stress contours between CSFEM and FEM at the ultimate bearing capacity.
6.4 Tunnel excavation
A tunnel excavation problem is considered in this example, with the geometry shown in Fig. 15(a). The tunnel has a diameter of 2 m and is excavated in five stages, each 0.4 m thick. The model is constrained by fixed boundaries at the bottom and normal constraints on the lateral sides. The surrounding rock is characterized by an elastic modulus of $E=8$ MPa, Poisson’s ratio $\nu=0.27$, cohesion $c=120$ kPa, and an internal friction angle of $40^{\circ}$. The analysis begins with an in-situ stress equilibrium step, followed by excavation simulation using the element birth–death technique.
Fig. 16 shows the vertical displacement at the tunnel crown during the different excavation stages, from which it can be observed that CSFEM and FEM exhibit good agreement. When the excavation reaches the fourth stage, the vertical displacement nearly converges, with a maximum value of –0.021 m. Moreover, Fig. 17 and Fig. 18 present the vertical displacement and stress contour plots at the 1st, 3rd, and 5th excavation stages, further demonstrating the close agreement between CSFEM and FEM.
Figure 15: The geometry and mesh model of tunnel excavation; (a) geometry model; (b) mesh model.
Figure 16: The vertical displacement of CSFEM and FEM under different excavation steps for the crown of the circular tunnel.
Figure 17: The vertical displacement of CSFEM and FEM under different excavation steps.
Figure 18: The vertical stress of CSFEM and FEM under different excavation steps.
6.5 Slope stability analysis
To verify the accuracy of the proposed method, we examined a uniform slope with a height of $H=10\,\text{m}$, sloping at an angle of $\beta=45^{\circ}$ with a friction angle of $\phi=20^{\circ}$, a unit weight of $\gamma=20\,\text{kN/m}^{3}$, and a cohesion of $c=12.38\,\text{kPa}$, as shown in Fig. 19. Based on the soil properties, the slope has a factor of safety of exactly 1.0, as per the limit analysis solution presented by Chen (Chen and Liu, 2012). The side and bottom boundaries of slopes are constrained without displacement $(\Delta X=0;\,\Delta Y=0)$ and $(\Delta X=0;\,\Delta Y=0;\,\Delta Z=0)$, respectively.
Figure 19: Geometry and boundary conditions of the slope
The stability of slope is evaluated using the shear strength reduction technique (Matsui and San, 1992). The parameters of shear strength are defined as
$c_{m}=\frac{c}{F_{r}}$ (35)
$\varphi_{m}=\arctan\biggl(\frac{\tan\varphi}{F_{r}}\biggr)$ (36)
where $c$ and $\varphi$ are the actual shear strength parameters, $c_{m}$ and $\varphi_{m}$ are the shear strength which can maintain the stability of slope. $F_{r}$ is the strength reduction factor. In the reduction process, the input shear strength parameters $c$ and $\varphi$ are reduction to $c_{m}$ and $\varphi_{m}$, to trigger the slope failure. The failure criterion was defined by the big jump in the nodal displacement at a chosen point close to the slope’s surface.
We extracted the horizontal displacement of the slope vertex and plotted the relationship between horizontal displacement and the safety factor, as shown in Fig. 20. A big jump in displacement occurred in CS-FEM when the safety factor reached 1.02, whereas in FEM, the displacement jump occurred at a safety factor of 0.99. It is worth noting that the safety factor obtained by CS-FEM is very close to the value of 1.0 reported in the literature  (Dawson et al., 1999) for this slope. Moreover, Fig. 21 shows the distribution of plastic strain in the slope at failure. It can be seen from the figure that the plastic zones have fully penetrated when the slope fails.
Figure 20: The relationship between displacement and the FOS; (a) the CSFEM; (b) the FEM.
Figure 21: The distribution of plastic strain in the slope at failure ; (a) the CSFEM; (b) the FEM.
7 Conclusions
This paper presented an elastic–plastic cell-based smoothed finite element method (EP-CSFEM) tailored for geotechnical engineering. Based on the theoretical formulation, numerical implementation, and extensive verification, the following key conclusions can be drawn:
Enhanced accuracy and stability – The incorporation of strain smoothing significantly improves stress recovery and reduces mesh dependency, effectively mitigating volumetric locking in nearly incompressible soils.
Robust elastic–plastic performance – The proposed return-mapping scheme ensures accurate stress updates and stable convergence in nonlinear problems, even under large deformations.
Versatile applicability – Benchmark tests and practical examples (pressurized cylinder, biaxial soil test, strip footing, tunnel excavation, and slope stability) confirm that EP-CSFEM matches or surpasses conventional FEM in predicting displacements, stresses, and ultimate bearing capacities.
Implementation efficiency – The ABAQUS UEL framework enables direct output of stress–strain fields, facilitating seamless integration with standard post-processing tools.
Overall, the CSFEM provides a reliable and efficient numerical platform for analyzing complex soil–structure interaction and slope stability problems. Future work will extend the method to three-dimensional large-deformation analyses and coupled hydro-mechanical processes.
8 Acknowledgements
The Yunnan Fundamental Research Projects (grant NO. 202401CF070043), the Xing Dian Talent Support Program of Yunnan Province (XDYC-QNRC-2022-0764) provided support for this study.